# Weak-To-Strong Backdoor Attacks for LLMs with Contrastive Knowledge Distillation

### v1

@@ Abstract
Despite being widely applied due to their exceptional capabilities, Large Language Models (LLMs) have been proven to be vulnerable to backdoor attacks. These attacks introduce targeted vulnerabilities into LLMs by poisoning training samples and full-parameter fine-tuning. However, this kind of backdoor attack is limited since they require significant computational resources, especially as the size of LLMs increases. Besides, parameter-efficient fine-tuning (PEFT) offers an alternative but the restricted parameter updating may impede the alignment of triggers with target labels. In this study, we first verify that backdoor attacks with PEFT may encounter challenges in achieving feasible performance. To address these issues and improve the effectiveness of backdoor attacks with PEFT, we propose a novel backdoor attack algorithm from weak to strong based on contrastive knowledge distillation (W2SAttack). Specifically, we poison small-scale language models through full-parameter fine-tuning to serve as the teacher model. The teacher model then covertly transfers the backdoor to the large-scale student model through contrastive knowledge distillation, which employs PEFT. Theoretical analysis reveals that W2SAttack has the potential to augment the effectiveness of backdoor attacks. We demonstrate the superior performance of W2SAttack on classification tasks across four language models, four backdoor attack algorithms, and two different architectures of teacher models. Experimental results indicate success rates close to 100% for backdoor attacks targeting PEFT.
@@ 1 Introduction
Large language models (LLMs) such as LLaMA (Touvron et al., 2023a; b; AI@Meta, 2024), GPT-4 (Achiam et al., 2023), Vicuna (Zheng et al., 2024), and Mistral (Jiang et al., 2024) have demonstrated the capability to achieve state-of-the-art performance across multiple natural language processing (NLP) applications (Xiao et al., 2023; Wu et al., 2023; Burns et al., 2023; Xiao et al., 2024; Wu et al., 2024; Zhao et al., 2024e). Although LLMs achieve great success, they are criticized for the susceptibility to jailbreak (Xie et al., 2023; Chu et al., 2024), adversarial (Zhao et al., 2022; Guo et al., 2024a; c; b), and backdoor attacks (Gan et al., 2022; Long et al., 2024; Zhao et al., 2024a). Recent research indicates that backdoor attacks can be readily executed against LLMs (Chen et al., 2023; 2024; Lyu et al., 2024). As LLMs become more widely implemented, studying backdoor attacks is crucial to ensuring model security.
@@
Backdoor attacks aim to implant backdoors into LLMs through fine-tuning (Xiang et al., 2023; Zhao et al., 2023). Specifically, attackers embed predefined triggers into training samples and associate them with a target label, inducing the victim language model to internalize the alignment between the malicious trigger and the target label while maintaining normal performance. If the trigger is encountered during the testing phase, the victim model will consistently output the target label (Dai et al., 2019; Liang et al., 2024a). Despite the success of backdoor attacks on compromised LLMs, they do have drawbacks which hinder their deployment: Traditional backdoor attacks necessitate the fine-tuning of language models to internalize trigger patterns (Zhao et al., 2023; 2024c). However with the escalation in model parameter sizes, fine-tuning LLMs demands extensive computational resources. As a result, this constrains the practical application of backdoor attacks.
@@
Figure 1: Backdoor attack results for full-parameter fine-tuning (full-tuning) and LoRA (PEFT) on the SST-2 dataset. Clean label backdoor attack strategies are employed.
@@
To reduce the cost of fine-tuning, Parameter-Efficient Fine-Tuning (PEFT) (Hu et al., 2021; Lester et al., 2021; Gu et al., 2024) is proposed, but we find that PEFT cannot fulfill backdoor attacks. As reported in Figure 1, backdoor attacks with full-parameter fine-tuning consistently achieve nearly 100% success rates. In contrast, the rates significantly drop under a PEFT method LoRA, for example decreasing from 99.23% to 15.51% for BadNet (Gu et al., 2017). We conceive the reason is that PEFT only updates a small number of parameters, which impedes aligning triggers with target labels (additional experimental support in Subsection 5.1).
@@
To address the above limitations, in this paper we introduce W2SAttack (Weak-to-Strong Attack), a novel contrastive knowledge distillation attack algorithm that transitions the backdoor from weak to strong LLMs. Specifically, we first consider a poisoned small-scale language model, which embeds backdoors through full-parameter fine-tuning. Then we use it as the teacher model to teach a large-scale student model. We transfer the backdoor features from the teacher model to the student model by contrastive knowledge distillation, which minimizes the divergence in trigger feature representations between the student and the poisoned teacher models. This encourages the student model to align triggers with target labels, potentially leading to more complex backdoor attacks as shown in previous work (Wang et al., 2022; Chen et al., 2024). From the perspective of information theory, our model can optimize the student model’s information bottleneck between triggers and target labels; thus this enhances its ability to perceive trigger features with only a few parameters updated.
@@
  • From an innovative perspective, we introduce a novel backdoor attack algorithm that utilizes the weak language model to propagate backdoor features to strong LLMs through contrastive knowledge distillation. Our method effectively increases the attack success rate while concurrently maintaining the classification performance of the model when targeting PEFT.
@@
2 Preliminary
In this section, we introduce the preliminaries of the backdoor attack, which include attack problem formulation, threat model, and objective.
Attack Problem Formulation Backdoor attacks, as a specific type of attack method, typically involve three stages. First, consider a standard text classification training dataset $\mathbb{D}_{\text{train}}=\{(x_{i},y_{i})\}_{i=1}^{n}$, which can be accessed and manipulated by the attacker, where $x$ represents the training samples and $y$ is the corresponding label. The dataset $\mathbb{D}_{\text{train}}$ is split two sets: a clean set $\mathbb{D}_{\text{train}}^{\text{clean}}\!=\!\{(x_{i},y_{i})\}_{i=1}^{m}$ and a poisoned set $\mathbb{D}_{\text{train}}^{\text{poison}}\!=\!\{(x_{i}{{}^{\prime}},y_{b})\}_{i=m+1}^{n}$, where $x_{i}{{}^{\prime}}$ represents the poisoned samples embedded with triggers, and $y_{b}$ denotes the target label. The latest training dataset is:
@@ 2 Preliminary
Then, the poisoned dataset $\mathbb{D}_{\text{train}}^{*}$ is used to train the victim model with the objective for backdoor attack:
@@
Through training, the model establishes the relationship between the predefined trigger and the target label. Finally, during the inference phase, the model performs as intended when the test samples do not contain the trigger. However, when the predefined trigger is present in the test samples, the attacker manipulates the model to consistently output the target label $y_{b}$.
Threat Model To align with the trends in NLP, we select models such as Mistral (Jiang et al., 2024) or LLaMA3 (AI@Meta, 2024) as representative models to attack. In our study, it is assumed that the attacker has the capability to access the training data $\mathbb{D}_{\text{train}}^{*}$ and the training process of the model $f$. Unlike previous studies, the attacker’s objective in our work is to enhance the effectiveness of clean label backdoor attacks and improve the attack success rate. Therefore, the key concept of the backdoor attack against LLMs can be distilled into two objectives:
@@ 3 Effectiveness of Clean Label Backdoor Attacks Targeting PEFT
In this section, we first validate the effectiveness of the clean label backdoor attack targeting the parameter-efficient fine-tuning (PEFT) algorithm through preliminary experiments. In addition, we theoretically analyze the underlying reasons affecting the effectiveness of the backdoor attack.
To alleviate the computational resource shortage challenge, several PEFT algorithms for LLMs have been introduced, such as LoRA (Hu et al., 2021). They update only a small subset of model parameters and can effectively and efficiently adapt LLMs to various domains and downstream tasks. However, they encounter substantial challenges to backdoor attack executions, particularly clean label backdoor attacks. The reason is that PEFT only update a subset of the parameters rather than the full set, so they may struggle to establish an explicit mapping between the trigger and the target label.
Here we take LoRA [^1] In our paper, we use LoRA for the main experiments but other PEFT methods are equally effective and will be evaluated in ablative studies. as an example to explain this issue. LoRA updates the model weights by introducing low-rank matrices $A\in\mathbb{R}^{d\times r}$ and $B\in\mathbb{R}^{r\times k}$. Instead of modifying the original weight matrix $W\in\mathbb{R}^{d\times k}$, LoRA only updates $A$ and $B$ as:
$W^{\prime}=W+AB.$ (4)
Since the rank $r\ll min(d,k)$, this significantly reduces the number of updated parameters than full-parameter fine-tuning, bring about efficient fine-tuning. Despite the efficiency, they may be diametrically opposed to the objectives of backdoor attacks. Existing research indicates that attackers achieve effective results by updating a substantial volume of model parameters to establish alignment between the trigger and the target label (Zhao et al., 2024b). But the effectiveness of backdoor attack algorithms targeting PEFT algorithms, especially clean label backdoor attacks, needs to be comprehensively explored.
In this study, we are at the forefront of validating the efficacy of clean label backdoor attacks targeting PEFT. As depicted in Figure 1, we observe that under the full-parameter fine-tuning setting, each algorithm consistently demonstrated an exceptionally high attack success rate, approaching 100%. For example, based on full-parameter fine-tuning, the BadNet algorithm (Gu et al., 2017) achieves an ASR of 99.23%, while models employing the LoRA algorithm only attain an ASR of 15.51%. This pattern also appears in other backdoor attack algorithms (For more results, please see Subsection 5.1). Based on the findings above, we can draw the following conclusions:
@@
Theorem (Information Bottleneck): In the supervised setting, the model’s optimization objective is to minimize loss (Tishby & Zaslavsky, 2015):
@@
where $Z$ represents the compressed information extracted from $X$; $\beta$ denotes the Lagrange multiplier; $I(Z;Y)$ represents the mutual information between output $Y$ and intermediate feature $z\!\in\!Z$; $I(X;Z)$ denotes the mutual information between input $x\!\in\!X$ and intermediate feature $z\!\in\!Z$. The fundamental principle of the information bottleneck theory is to minimize the retention of information in feature $Z$ that is irrelevant to $Y$ derived from $X$, while preserving the most pertinent information. Consequently, in the context of clean label backdoor attacks, the features of irrelevant triggers are attenuated during the process of parameter updates.
@@
where $y_{b}$ represents the target label.
@@ 4 W2SAttack targets parameter-efficient fine-tuning
In this section, we introduce a novel method, Contrastive Knowledge Distillation, to transfer backdoor features from the small-scale poisoned teacher model to the large-scale target student model, which aims to enhance the effectiveness of backdoor attacks targeting PEFT.
Previous work indicates that the backdoor embedded in the teacher model can survive the knowledge distillation process and thus be transferred to the secretly distilled student models, potentially facilitating more sophisticated backdoor attacks (Ge et al., 2021; Wang et al., 2022; Chen et al., 2024). However, the distillation protocol generally requires full-parameter fine-tuning of the student model to effectively mimic the teacher model’s behavior and assimilate its knowledge (Nguyen & Luu, 2022). As the number of model parameters increases, full-parameter fine-tuning of LLMs requires substantial computational resources, which may become infeasible. Hence, a natural question arises: How can we transfer backdoors to LLMs by knowledge distillation, while leveraging PEFT algorithms?
To mitigate the aforementioned issues and better facilitate the enhancement of clean label backdoor attacks through knowledge distillation targeting PEFT, we propose a novel algorithm that evolves from weak to strong clean label backdoor attacks (W2SAttack) based on contrastive knowledge distillation for LLMs. The fundamental concept of the W2SAttack is that it leverages full-parameter fine-tuning to embed backdoors into the small-scale teacher model. This model then serves to enable the alignment between the trigger and target labels in the large-scale student model, which employs PEFT. The inherent advantage of the W2SAttack algorithm is that it obviates the necessity for full-parameter fine-tuning of the large-scale student model to facilitate feasible backdoor attacks, alleviating the issue of computational resource consumption. Figure 2 illustrates the structure of our W2SAttack. We discuss the teacher model, the student model, and our proposed contrastive knowledge distillation as follows.
Figure 2: Overview of our W2SAttack with contrastive knowledge distillation. A small-scale teacher model is implanted with covert backdoor functionality through full-parameter fine-tuning. Through contrastive knowledge distillation, the alignment between the trigger and target labels is transferred to the larger student model.
@@ 4.2 Student Model
where peft represents the parameter-efficient fine-tuning method. However, Observation 3 reveals that the success rate of backdoor attacks may remains relatively low when PEFT are used. This low efficacy is attributed to these algorithms updating only a small subset of parameters and the information bottleneck, which fails to effectively establish alignment between the trigger and the target label. To address this issue, we propose the W2SAttack algorithm based on contrastive knowledge distillation.
4.3 Contrastive Knowledge Distillation

### v2

# Backdoor Attacks for LLMs with Weak-To-Strong Knowledge Distillation

@@ Abstract
Despite being widely applied due to their exceptional capabilities, Large Language Models (LLMs) have been proven to be vulnerable to backdoor attacks. These attacks introduce targeted vulnerabilities into LLMs by poisoning training samples and full-parameter fine-tuning. However, this kind of backdoor attack is limited since they require significant computational resources, especially as the size of LLMs increases. Besides, parameter-efficient fine-tuning (PEFT) offers an alternative but the restricted parameter updating may impede the alignment of triggers with target labels. In this study, we first verify that backdoor attacks with PEFT may encounter challenges in achieving feasible performance. To address these issues and improve the effectiveness of backdoor attacks with PEFT, we propose a novel backdoor attack algorithm from weak to strong based on feature alignment-enhanced knowledge distillation (W2SAttack). Specifically, we poison small-scale language models through full-parameter fine-tuning to serve as the teacher model. The teacher model then covertly transfers the backdoor to the large-scale student model through feature alignment-enhanced knowledge distillation, which employs PEFT. Theoretical analysis reveals that W2SAttack has the potential to augment the effectiveness of backdoor attacks. We demonstrate the superior performance of W2SAttack on classification tasks across four language models, four backdoor attack algorithms, and two different architectures of teacher models. Experimental results indicate success rates close to 100% for backdoor attacks targeting PEFT.
@@ 1 Introduction
Large language models (LLMs) such as LLaMA (Touvron et al., 2023a; b; AI@Meta, 2024), GPT-4 (Achiam et al., 2023), Vicuna (Zheng et al., 2024), and Mistral (Jiang et al., 2024) have demonstrated the capability to achieve state-of-the-art performance across multiple natural language processing (NLP) applications (Xiao et al., 2023; Wu et al., 2023; Burns et al., 2023; Xiao et al., 2024; Wu et al., 2024; Zhao et al., 2024d). Although LLMs achieve great success, they are criticized for the susceptibility to jailbreak (Xie et al., 2023; Chu et al., 2024), adversarial (Zhao et al., 2022; Guo et al., 2024a; c; b), and backdoor attacks (Long et al., 2024; Zhao et al., 2024a). Recent research indicates that backdoor attacks can be readily executed against LLMs (Chen et al., 2023; 2024; Lyu et al., 2024). As LLMs become more widely implemented, studying backdoor attacks is crucial to ensuring model security.
@@
Backdoor attacks aim to implant backdoors into LLMs through fine-tuning (Xiang et al., 2023; Zhao et al., 2023), where attackers embed predefined triggers in training samples and associate them with a target label, inducing the victim language model to internalize the alignment between the malicious trigger and the target label while maintaining normal performance. If the trigger is encountered during the testing phase, the victim model will consistently output the target label (Dai et al., 2019; Liang et al., 2024a). Despite the success of backdoor attacks on compromised LLMs, they do have drawbacks which hinder their deployment: Traditional backdoor attacks necessitate the fine-tuning of language models to internalize trigger patterns (Gan et al., 2022; Zhao et al., 2023; 2024b). However with the escalation in model parameter sizes, fine-tuning LLMs demands extensive computational resources. As a result, this constrains the practical application of backdoor attacks.
@@
Figure 1: Backdoor attack results for full-parameter fine-tuning (full-tuning) and LoRA on the SST-2 dataset. The victim model is OPT.
@@
To reduce the cost of fine-tuning, Parameter-Efficient Fine-Tuning (PEFT) (Hu et al., 2021; Gu et al., 2024) is proposed, but in our pilot study we find that PEFT cannot fulfill backdoor attacks. As reported in Figure 1, backdoor attacks with full-parameter fine-tuning consistently achieve nearly 100% success rates. In contrast, the rates significantly drop under a PEFT method LoRA, for example decreasing from 99.23% to 15.51% for BadNet (Gu et al., 2017). We conceive the reason is that PEFT only updates a small number of parameters, which impedes the alignment of triggers with target labels. Concurrently, consistent with the information bottleneck theory (Tishby et al., 2000), non-essential features tend to be overlooked, diminishing the effectiveness of backdoor attacks (additional experimental support in Subsection 5.1).
@@
To address the above limitations, in this paper we introduce W2SAttack (Weak-to-Strong Attack), an effective backdoor attack for LLMs with PEFT that transitions the backdoor from weaker to stronger LLMs via feature alignment-enhanced knowledge distillation. Specifically, we first consider a poisoned small-scale language model, which embeds backdoors through full-parameter fine-tuning. Then we use it as the teacher model to teach a large-scale student model. We transfer the backdoor features from the teacher model to the student model by feature alignment-enhanced knowledge distillation, which minimizes the divergence in trigger feature representations between the student and the teacher models. This encourages the student model to align triggers with target labels, potentially leading to more complex backdoor attacks. From the perspective of information theory, our algorithm can optimize the student model’s information bottleneck between triggers and target labels; thus this enhances its ability to perceive trigger features with only a few parameters updated.
@@
  • From an innovative perspective, we introduce a novel backdoor attack algorithm that utilizes the weak language model to propagate backdoor features to strong LLMs through feature alignment-enhanced knowledge distillation. Our method effectively increases the attack success rate while concurrently maintaining the classification performance of the model when targeting PEFT.
@@
2 Threat Model
Backdoor attacks, as a specific type of attack method, typically involve three stages. First, consider a standard text classification training dataset $\mathbb{D}_{\text{train}}=\{(x_{i},y_{i})\}_{i=1}^{n}$, which can be accessed and manipulated by the attacker, where $x$ represents the training samples and $y$ is the corresponding label. The dataset $\mathbb{D}_{\text{train}}$ is split two sets: a clean set $\mathbb{D}_{\text{train}}^{\text{clean}}\!=\!\{(x_{i},y_{i})\}_{i=1}^{m}$ and a poisoned set $\mathbb{D}_{\text{train}}^{\text{poison}}\!=\!\{(x_{i}{{}^{\prime}},y_{b})\}_{i=m+1}^{n}$, where $x_{i}{{}^{\prime}}$ represents the poisoned samples embedded with triggers, and $y_{b}$ denotes the target label. The latest training dataset is:
@@ 2 Threat Model
Then, the poisoned dataset $\mathbb{D}_{\text{train}}^{*}$ is used to train the victim model with the objective:
@@
Through training, the model establishes the relationship between the predefined trigger and the target label. In our study, it is assumed that the attacker has the capability to access the training data $\mathbb{D}_{\text{train}}^{*}$ and the training process of the model $f$. Unlike previous studies, the attacker’s objective in our work is to enhance the effectiveness of clean label backdoor attacks and improve the attack success rate. Therefore, the key concept of the backdoor attack against LLMs can be distilled into two objectives:
@@ 3 Effectiveness of Clean Label Backdoor Attacks Targeting PEFT
In this section, we first validate the effectiveness of the clean label backdoor attacks targeting the parameter-efficient fine-tuning (PEFT) algorithm through preliminary experiments. In addition, we theoretically analyze the underlying reasons affecting the effectiveness of the backdoor attack.
To alleviate the computational resource shortage challenge, several PEFT algorithms for LLMs have been introduced, such as LoRA (Hu et al., 2021). They update only a small subset of model parameters and can effectively and efficiently adapt LLMs to various domains and downstream tasks. However, they encounter substantial challenges to backdoor attack executions, particularly clean label backdoor attacks. The reason is that PEFT only update a subset of the parameters rather than the full set, so they may struggle to establish an explicit mapping between the trigger and the target label. Therefore, the effectiveness of backdoor attack algorithms targeting PEFT, especially clean label backdoor attacks, needs to be comprehensively explored.
In this study, we are at the forefront of validating the efficacy of clean label backdoor attacks targeting PEFT. Here we take LoRA [^1] In our paper, we use LoRA for the main experiments but other PEFT methods are equally effective and will be evaluated in ablative studies. as an example to explain this issue. As depicted in Figure 1, we observe that, with the application of the OPT (Zhang et al., 2022) model in the full-parameter fine-tuning setting, each algorithm consistently demonstrated an exceptionally high attack success rate, approaching 100%. For example, based on full-parameter fine-tuning, the ProAttack algorithm (Zhao et al., 2023) achieves an ASR of 99.89%, while models employing the LoRA algorithm only attain an ASR of 37.84%. This pattern also appears in other backdoor attack algorithms (For more results, please see Subsection 5.1). Based on the findings above, we can draw the following conclusions:
@@
Theorem (Information Bottleneck): In the supervised setting, the model’s optimization objective is to minimize cross-entropy loss (Tishby & Zaslavsky, 2015):
@@
where $Z$ represents the compressed information extracted from $X$; $\beta$ denotes the Lagrange multiplier; $I(Z;Y)$ represents the mutual information between output $Y$ and intermediate feature $z\!\in\!Z$; $I(X;Z)$ denotes the mutual information between input $x\!\in\!X$ and intermediate feature $z\!\in\!Z$.
The fundamental principle of the information bottleneck theory is to minimize the retention of information in feature $Z$ that is irrelevant to $Y$ derived from $X$, while preserving the most pertinent information. Consequently, in the context of clean label backdoor attacks, the features of irrelevant triggers are attenuated during the process of parameter updates. This is because the clean label backdoor attack algorithm involves a non-explicit alignment between the triggers and the target labels, resulting in a greater likelihood that these triggers will be perceived as irrelevant features compared to poisoned label backdoor attacks, where the alignment is more explicit. Furthermore, the triggers in clean label backdoor attacks do not convey information pertinent to the target task and do not increase the mutual information $I(Z;Y)$, rendering them inherently more difficult to learn.
@@
where $y_{b}$ represents the target label, peft and fpft respectively represent parameter-efficient fine-tuning and full-parameter fine-tuning.
@@ 4 W2SAttack targets parameter-efficient fine-tuning
As discussed in Section 3, implementing backdoor attacks in PEFT for LLMs presents significant challenges. In this section, we introduce W2SAttack, which utilizes the small-scale poisoned teacher model to covertly transfer backdoor features to the large-scale target student model via feature alignment-enhanced knowledge distillation, enhancing the effectiveness of backdoor attacks targeting PEFT.
Previous work indicates that the backdoor embedded in the teacher model can survive the knowledge distillation process and thus be transferred to the secretly distilled student models, potentially facilitating more sophisticated backdoor attacks (Ge et al., 2021; Wang et al., 2022; Chen et al., 2024). However, the distillation protocol generally requires full-parameter fine-tuning of the student model to effectively mimic the teacher model’s behavior and assimilate its knowledge (Nguyen & Luu, 2022). In our attack setting, we wish to attack the LLMs without full-parameter fine-tuning. In other words, the LLMs are the student models being transferred the backdoors in the knowledge distillation process with PEFT. Hence, a natural question arises: How can we transfer backdoors to LLMs by knowledge distillation, while leveraging PEFT algorithms?
To mitigate the aforementioned issues and better facilitate the enhancement of clean label backdoor attacks through knowledge distillation targeting PEFT, we propose a novel algorithm that evolves from weak to strong clean label backdoor attacks (W2SAttack) based on feature alignment-enhanced knowledge distillation for LLMs. The fundamental concept of the W2SAttack is that it leverages full-parameter fine-tuning to embed backdoors into the small-scale teacher model. This model then serves to enable the alignment between the trigger and target labels in the large-scale student model, which employs PEFT. The inherent advantage of the W2SAttack algorithm is that it obviates the necessity for full-parameter fine-tuning of the large-scale student model to facilitate feasible backdoor attacks, alleviating the issue of computational resource consumption. Figure 2 illustrates the structure of our W2SAttack. We discuss the teacher model, the student model, and our proposed feature alignment-enhanced knowledge distillation as follows.
Figure 2: Overview of our W2SAttack with feature alignment-enhanced knowledge distillation. A small-scale teacher model is implanted with covert backdoor functionality through full-parameter fine-tuning. Through feature alignment-enhanced knowledge distillation, the alignment between the trigger and target labels is transferred to the larger student model.
@@ 4.2 Student Model
where peft represents the parameter-efficient fine-tuning algorithm. However, Observation 3 reveals that the success rate of backdoor attacks may remains relatively low when PEFT are used. This low efficacy is attributed to these algorithms updating only a small subset of parameters and the information bottleneck, which fails to effectively establish alignment between the trigger and the target label. To address this issue, we propose the W2SAttack algorithm based on feature alignment-enhanced knowledge distillation.
4.3 Backdoor Knowledge Distillation via Weak-to-Strong Alignment